how would you solve this word problem?
Bob has a garden that is 9 feet long by 6 feet wide. He wants to increase each dimension of the garden by the same amount and have the area of the new garden be 5 times the area of the present garden. What should the new dimensions be?

Answer by richwmiller(17219)

5*(9*6)=(x+9)*(x+6)
x=9
check
5*(9*6)=18*15
5*3*3*6=18*15
ok